OpenStudy (anonymous):

A student studying for a vocabulary test knows the meanings of 12 words from a list of 24 words. If the test contains 10 words from the study list, what is the probability that at least 8 of the words on the test are words that the student knows? (Round your answer to three decimal places.)

OpenStudy (anonymous):

P(8) + P(9) + P(10)

OpenStudy (anonymous):

@sourwing How do I calculate this? I always mess these up

OpenStudy (anonymous):

O_O...

OpenStudy (anonymous):

IDK how to do this.

OpenStudy (anonymous):

for \(P(x=8)\) the denominator will be the number of ways to choose 10 out of the 24 questions the numerator will be the number of ways to choose 8 out of the 12 known, times the number of ways to choose 2 out of the 24 unknown

OpenStudy (anonymous):

i had a typo above, i should have written \[the numerator will be the number of ways to choose 8 out of the 12 known, times the number of ways to choose 2 out of the 12 unknown

OpenStudy (anonymous):

whole this is a typo i must be tired let me try again

OpenStudy (anonymous):

\[P(x=8)=\frac{\binom{12}{8}\times \binom{12}{2}}{\binom{24}{10}}\]

OpenStudy (anonymous):

that is better

OpenStudy (anonymous):

\[P(x=9)=\frac{\binom{12}{9}\times \binom{12}{1}}{\binom{24}{10}}\]

OpenStudy (anonymous):

Thank you!